# Task-driven SLAM Benchmarking for Robot Navigation

### v2

@@ III-C Open-Loop EuRoC Evaluation - A Preliminary Study
This section describes the performance evaluation performed for the provisional SLAM performance analysis noted in the Introduction, which argued that performance evaluation would benefit from including task-driven evaluation. The EuRoC machine hall sequences may be interpreted as an aerial vehicle inspection task; the objective is to consistently revisit specific points during repeat inspections. Evaluation involves pose estimation accuracy and precision, with each frame treated as an inspection point to reach.
@@ III-C2 Results and Analysis
Fig. 1 visualizes the results as Precision vs. Accuracy plots. Given that the red dashed line is precision-accuracy parity, position precision values match or are lower than accuracy values across all sequences. Orientation precision is noticeably better than accuracy. The results exhibit better localization repeatability than absolute accuracy would suggest. When utilizing a SLAM map, both methods exhibit a improved performance with outcomes shifted toward the lower-left corner. Position precision error, which is within $1/6$ of a robot diameter in w/o map tests, reduces to $1/12$ of a robot diameter in w/map tests. Orientation precision error, at $1/35\,FoV$, reduces by a factor of $10$ in the w/map tests (to pixel level error). An object intended to be centered in the field of view from a reasonable distance would still be visible and nearly centered. This suggests that accuracy does not fully reflect potential SLAM performance from a task-driven perspective, and points towards the benefits of augmenting it with precision.
@@ III-D1 Simulation
Fig. 5 Precision vs Accuracy plots in without map tests for position and orientation consistently show lower precision value than accuracy when the navigation succeeds. To some degree this indicates that localization failure is catastrophic when it occurs; there doesn’t seem to be a sliding scale. The cumulative position precision plot in Fig. 6 indicates the same by plateauing rather than exhibiting linear growth to 100% across the precision axis. In the cases studied, either a SLAM method is precise to a robot diameter or it fails. Orientation accuracy and precision values are low given the field of view. Due to the strong performance of orientation, it will not be reviewed in future analyses.
The least complete implementations are odometry methods (DSOL, SVO-Pro), MSCKF and as they cannot leverage map content, nor long baseline associations outside of the filtering window. The two 3D-LiDAR methods performed comparably in the w/o map tests. However, they did not match 2D-LiDAR performance. This performance contrasts with their near universal preference in robot deployments [23], which may be due to the heavy emphasis on exploration rather than repeatable navigation. Exploration has looser navigation demands. HectorSLAM and ORB_SLAM3 were closer to perfect, but exhibit enough failures to not be reliable. One visual SLAM method, GF-GG, matched the perfect completeness performance of the 2D-LiDAR SLAM-Toolbox implementation. Both achieved 100% cumulative precision at one robot diameter, indicating that the two methods are comparable in performance, though SLAM-Toolbox has an edge in the N-AUC score.
Using a map improves performance as seen in Fig. 6 by the upwards and/or leftwards shift of plotted curves relative to the w/o map curve. Correspondingly, the N-AUC scores improve. SLAM-Toolbox shows minimal change, primarily due to its scan-to-map registration design, which is well-suited for indoor scenarios, and its continuous pose-graph-based global map management. ORB_SLAM3’s imperfect w/o map completion becomes perfect w/map completion. It has the lowest N-AUC score relative to SLAM-Toolbox and GF-GG, and hits perfect completion at around $1.5D$.
@@ III-D2 Real-World
Analysis will focus only on position precision, as the average orientation precision value lies below 10 degrees for all methods. Being at $1/8$ of the camera’s field of view, this level of orientation precision is acceptable from a task-driven perspective. Fig. 8 provides cumulative position precision plots for the real-world tests (w/o and w/map). HectorSLAM, DSOL, and MSCKF did not perform well enough to warrant inclusion (below 20%). Performance of the other implementations roughly matches the ordering of simulation, with SLAM-Toolbox and GF-GG achieving perfect completion but at around 2 robot diameters, which may influence the robot’s ability to perceive the task workspace region upon arrival. The LIO methods and ORB_SLAM3 achieved high completion rates. Passive stereo implementations continue to exhibit comparable performance to LiDAR-based implementations.
Moving to the Fig. 8 w/map case, there is a left-ward shift in the curve indicating improved precision and an overall compression in pose variance across methods. The violin plots of Fig. 8 (top), support this same finding. Only SLAM-Toolbox manages $100\%$ completion, with the others failing on the second long distance test. The stereo SLAM methods failures are programmatic (program crashes), arising from issues managing the visual map over long time-and large spatial-scales, especially when there are periods of poor visual conditions. They achieved perfect performance on the short run test; for those runs GF-GG and ORB_SLAM3 matched SLAM-Toolbox with an N-AUC of 0.93 and 0.95 vs 0.92. The violin plots of GF-GG and ORB_SLAM3 consist of only the short run test. The degraded long-term performance indicates that future improvements should be aimed at back-end elements (long-term mapping, graph maintenance, and keyframe culling). Resolving these bugs may lead to performance matching that of SLAM-Toolbox. The multimedia attachment provides visual evidence from the cameras to qualitatively see how precision varies across the methods.
Refer to captionRefer to captionRefer to caption
@@ III-D3 Discussion
The entirety of TaskSLAM-Bench results reveal that Visual SLAM achieves performance approaching that of LiDAR SLAM in closed-loop navigation tasks, a finding not captured by existing benchmarks. These results demonstrate that vision-based approaches are capable of delivering strong navigation performance once robust program execution is achieved. Using a SLAM map with map-to-frame tracking significantly improves performance. It reduces drift and enhances long-term consistency, which has positive impacts on maintaining robust and accurate navigation.
@@
Overall, GF-GG outperforms ORB_SLAM3 by prioritizing efficient map-to-frame tracking and local BA, suggesting scalability improvements are critical to long-term SLAM. Better integrating robot odometry into the pose graph design can enhance robustness in degenerate, feature poor scenarios to ensure continuous operation.

### v3

@@ III-C Open-Loop EuRoC Evaluation - A Preliminary Study
This section describes the performance evaluation performed for the provisional SLAM performance analysis noted in the Introduction, which argued that performance evaluation would benefit from including task-driven evaluation. The commonly used EuRoC machine hall sequences ground this preliminary study to other SLAM benchmarking works. Given the lack of distinct waypoints, precision and accuracy are based on the pose estimates at each frame.
@@ III-C2 Results and Analysis
Fig. 1 visualizes the results as Precision vs. Accuracy plots. Given that the red dashed line is precision-accuracy parity, position precision values match or are lower than accuracy values across all sequences. Orientation precision is noticeably better than accuracy. The results exhibit better localization repeatability than absolute accuracy would suggest. When utilizing a SLAM map, both methods exhibit improved performance with outcomes shifted toward the lower-left corner. Position precision error, which is within $1/6$ of a robot diameter in w/o map tests, reduces to $1/12$ of a robot diameter in w/map tests. Orientation precision error, at 1/35FoV1/35\,FoV, reduces by a factor of $10$ in the w/map tests (to pixel level error). An object intended to be centered in the field of view from a reasonable distance would still be visible and nearly centered. This suggests that accuracy does not fully reflect potential SLAM performance from a task-driven perspective, and points towards the benefits of augmenting it with precision.
@@ III-D1 Simulation
Fig. 5 Precision vs Accuracy plots in without map tests consistently show lower precision value than accuracy when the navigation succeeds. To some degree this indicates that localization failure is catastrophic when it occurs; there doesn’t seem to be a sliding scale. The cumulative position precision plot in Fig. 6 indicates the same by plateauing rather than exhibiting linear growth to 100% across the precision axis. In the cases studied, either a SLAM method is precise to a robot diameter or it fails. Orientation accuracy and precision values are low given the field of view. Due to the strong performance of orientation, it will not be reviewed in future analyses.
The least complete implementations are odometry methods (DSOL, SVO-Pro), and MSCKF as they cannot leverage map content, nor long baseline associations outside of the filtering window. The two 3D-LiDAR methods performed comparably in the w/o map tests. However, they did not match 2D-LiDAR performance. This performance contrasts with their near universal preference in robot deployments [25], which may be due to the heavy emphasis on exploration rather than repeatable navigation. Exploration has looser navigation demands. HectorSLAM and ORB_SLAM3 were closer to perfect, but exhibit enough failures to not be reliable. One visual SLAM method, GF-GG, matched the perfect completeness performance of the 2D-LiDAR SLAM-Toolbox implementation. Both achieved 100% cumulative precision at one robot diameter, indicating that the two methods are comparable in performance, though SLAM-Toolbox has an edge in the N-AUC score.
Using a map improves performance as seen in Fig. 6 by the upwards and/or leftwards shift of plotted curves relative to the w/o map curve. Correspondingly, the N-AUC scores improve. SLAM-Toolbox shows minimal change, primarily due to its scan-to-map registration design, which is well-suited for indoor scenarios, and its continuous pose-graph-based global map management. ORB_SLAM3’s imperfect w/o map completion becomes perfect w/map completion. It has a lower N-AUC score relative to SLAM-Toolbox and GF-GG, hitting perfect completion at around $1.5D$.
@@ III-D2 Real-World
Analysis will focus only on position precision, as the average orientation precision value lies below 10 degrees for all methods. Being at $1/8$ of the camera’s field of view, this level of orientation precision is acceptable from a task-driven perspective. Fig. 8 provides cumulative position precision plots for the real-world tests (w/o and w/map). HectorSLAM, DSOL, and MSCKF did not perform well enough to warrant inclusion (below 35%). Performance of the other implementations roughly matches the ordering of simulation, with SLAM-Toolbox and GF-GG achieving perfect completion at one robot diameter. The LIO methods and ORB_SLAM3 achieved high completion rates. Passive stereo implementations continue to exhibit comparable performance to LiDAR-based implementations.
Looking at the Fig. 8 w/-map case, there is a left-ward shift in the curve indicating improved precision and an overall compression in pose variance across methods. The violin plots of Fig. 8 (top), support this same finding. SLAM-Toolbox and GF-GG manage $100\%$ completion (N-AUC: 0.93 vs. 0.90). ORB_SLAM3 and LIO-SAM show a substantial decrease in completion and N-AUC. This decrease is due to failure during the initial mapping phase on the long-distance test. In the short-distance run, GF-GG and ORB_SLAM3 performed comparably to SLAM-Toolbox and LIO-SAM, achieving N-AUC scores of 0.96 and 0.97 versus 0.94 and 0.86, respectively. GF-GG performs similarly across both testing modes, indicating that future improvements should target long-term mapping graph maintenance and map exploitation for localization in Visual SLAM. The multimedia attachment provides visual evidence from the cameras to qualitatively see how precision varies across methods.
Refer to captionRefer to captionRefer to captionw/o map Refer to captionw/ map Refer to caption
@@ III-D3 Discussion
The TaskSLAM-Bench results reveal that Visual SLAM achieves performance approaching that of LiDAR SLAM for closed-loop navigation tasks, a finding not captured by existing benchmarks. These results demonstrate that vision-based approaches are capable of delivering strong navigation performance once robust program execution is achieved. Using a SLAM map with map-to-frame tracking significantly improves performance. It reduces drift and enhances long-term consistency, which has positive impacts on maintaining robust and accurate navigation.
@@
Overall, GF-GG outperforms ORB_SLAM3 by prioritizing efficient map-to-frame tracking and local BA, suggesting scalability improvements are critical to long-term SLAM. Better integrating robot odometry into the pose graph design can enhance robustness during feature poor scenarios to ensure continuous operation.